OpenStudy (anonymous):

3(3 + 4x) - 7 is greater than or equal to 62 Solution set is {x|x is greater than or equal to ?|

OpenStudy (katrinakaif):

X is greater than or equal to 5.

OpenStudy (katrinakaif):

Why ? First add 7 on both sides.. which gives 62 + 7 = 69. then divide by 3 orn both sides giving 69/3 = 23. then substract by 3 on both sides giving 23-3 = 20 then divide by 4 on both sides giving 20/4 = 5 x is greathern than or equal to 5.